OpenStudy (anonymous):

I need to convert 4/5-6costheta to a rectangular equation

OpenStudy (anonymous):

I feel like I should start off with multiplying both sides by r but that gives me r in the numerator :/

OpenStudy (anonymous):

so your equation is r=4/5 - 6 cos theta and you want to convert it to rectangular form

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I've tried working it a few different ways and I always hit a dead end

OpenStudy (anonymous):

in rectangular form you aren't going to get a function

OpenStudy (anonymous):

I looked at a similar one, r= 6/7-8costheta and it gave the answer 15x^2+96x-49y^2+36=0 but I don't know how to get to that

OpenStudy (anonymous):

those two aren't the same at all, graph them both and see

OpenStudy (anonymous):

that's what my homework says.

OpenStudy (anonymous):

https://www.desmos.com/calculator/gyz60mxqdi

OpenStudy (anonymous):

you didn't enter it corrently

OpenStudy (anonymous):

I guess that's my fault 6/(7-8costheta)

OpenStudy (anonymous):

and the same with the other one?

OpenStudy (anonymous):

Yes, sorry :(

OpenStudy (anonymous):

so you can do it by analysis or by calculation. analysis is much easier, but won't get you very far if your teacher says to show your work

OpenStudy (anonymous):

the only example he gave for this type of problem was one much more simple: r=3costheta. He multiplied both sides of the equation by r, then changed the left to x^2+y^2 and changed 3rcos(theta) to 3x and just solved from there

OpenStudy (anonymous):

\[r=\frac{4}{5-6\cos \theta }\] \[r^2=\frac{4r}{5-6\cos \theta }\] \[x^2+y^2=\frac{4(x^2+y^2)}{5-6x }\]

OpenStudy (anonymous):

correction: 4(sqrt(x^2+y^2)) in the numerators

OpenStudy (anonymous):

okay, I was a little confused there. How did you change 6cos(theta) to 6x? I thought it had to be 6rcos(theta)?

OpenStudy (anonymous):

 sin theta = x/1, cos theta = y/1

OpenStudy (anonymous):

*sin theta = y/1, cos theta = x/1

OpenStudy (anonymous):

we can assume that r=1?

OpenStudy (anonymous):

for the unit circle, yes

OpenStudy (anonymous):

yes, but for the problem?

OpenStudy (anonymous):

in my notes and the video I watched on youtube it says cos(theta)=x/r

OpenStudy (anonymous):

so correction: \[x^2+y^2=\frac{4r}{5-6\frac{ x }{ r } }\] multiply by r/r \[x^2+y^2=\frac{4r^2}{5r-6x}\] multiply both sides by 5r-6x \[(5r-6)\times(x^2+y^2)=4x^2+4y^2\]

OpenStudy (anonymous):

hmmm, I went from there and divided each side by 4x^2+4y^2, factored out the 4 which made it cancel with the top. But again, it doesn't seem right

zepdrix (zepdrix):

Hmm you guys are approaching this a little weird :O I would probably do it like this:\[\Large r=\frac{4}{5-6\cos \theta}\]Multiply through by that denominator:\[\Large r(5-6\cos \theta)=4\]Distribute:\[\Large 5\color{orangered}{r}-6\color{royalblue}{r \cos \theta}=4\] Use your identities:\[\Large \color{royalblue}{r \cos \theta=x}\]\[\Large \color{orangered}{r=x^2+y^2}\] \[\Large 5\color{orangered}{(x^2+y^2)}-6\color{royalblue}{x}=4\]And then from there I guess you just need to complete the square on the x's to solve for y.

OpenStudy (anonymous):

Yeahhhh, I didn't trust what he was doing and was waiting for him to admit that he didn't know either, haha

zepdrix (zepdrix):

lol :)

OpenStudy (anonymous):

but I thought r=(x^2+y^2)^(1/2)?

zepdrix (zepdrix):

\[\Large \sin^2\theta+\cos^2\theta=1\]Multiply each side by r^2,\[\Large (\color{#DD4747 }{r \sin \theta})^2+(\color{royalblue}{r \cos \theta})^2=r^2\]From here we can plug in our identities:\[\Large \color{#DD4747 }{r \sin \theta=y}\]\[\Large \color{royalblue}{r \cos \theta =x}\] \[\Large (\color{#DD4747 }{y})^2+(\color{royalblue}{x})^2=r^2\]

zepdrix (zepdrix):

It's the form of a circle. It's how we relate trig back to rectangular, with circle stuff :p

OpenStudy (anonymous):

but we just had r, not r^2

zepdrix (zepdrix):

Ohhhh we just had r, XDDD haha good call.. I made a boo boo didn't I?

OpenStudy (anonymous):

:)

zepdrix (zepdrix):

So that last line should be:\[\Large 5\color{orangered}{\sqrt{x^2+y^2}}-6\color{royalblue}{x}=4\]

zepdrix (zepdrix):

Oh it wasn't a squared r.. doh! Then maybe pete's method was on the right track :3

OpenStudy (anonymous):

yeah, that's what I got when I continued his, yours was just a lot faster

zepdrix (zepdrix):

Hmm so if you want to solve for y, you have to do some messy steps from here I guess.\[\Large 5\sqrt{x^2+y^2}\quad=\quad 6x+4\]Squaring:\[\Large 25x^2+25y^2=36x^2+48x+16\]And then you can probably solve for y now :O

OpenStudy (anonymous):

ohhh, okay, I was making it a lot more messy by squaring the whole left side without moving the 6x to the right, haha

zepdrix (zepdrix):

Ooo yah that could be a little tricky.

OpenStudy (anonymous):

thank you :)

zepdrix (zepdrix):

np \c:/